# Truth Knows No Language: Evaluating Truthfulness Beyond English

### v1

@@ Abstract
We introduce a professionally translated extension of the TruthfulQA benchmark designed to evaluate truthfulness in Basque, Catalan, Galician, and Spanish. Truthfulness evaluations of large language models (LLMs) have primarily been conducted in English. However, the ability of LLMs to maintain truthfulness across languages remains under-explored. Our study evaluates 12 state-of-the-art open LLMs, comparing base and instruction-tuned models using human evaluation, multiple-choice metrics, and LLM-as-a-Judge scoring. Our findings reveal that, while LLMs perform best in English and worst in Basque (the lowest-resourced language), overall truthfulness discrepancies across languages are smaller than anticipated. Furthermore, we show that LLM-as-a-Judge correlates more closely with human judgments than multiple-choice metrics, and that informativeness plays a critical role in truthfulness assessment. Our results also indicate that machine translation provides a viable approach for extending truthfulness benchmarks to additional languages, offering a scalable alternative to professional translation. Finally, we observe that universal knowledge questions are better handled across languages than context- and time-dependent ones, highlighting the need for truthfulness evaluations that account for cultural and temporal variability. Dataset and code are publicly available under open licenses. [^1] https://github.com/hitz-zentroa/truthfulqa-multi and https://huggingface.co/datasets/HiTZ/truthfulqa-multi
@@
Blanca Calvo Figueras , Eneko Sagarzazu, Julen Etxaniz, Jeremy Barnes, Pablo Gamallo, Iria De Dios Flores Φ, Rodrigo Agerri HiTZ Center - Ixa, University of the Basque Country, UPV/EHU  Elhuyar Centro de Investigación en Tecnoloxías Intelixentes (CiTIUS), Universidade de Santiago de Compostela ΦDepartament de Traducció i Ciències del Llenguatge, Universitat Pompeu Fabra blanca.calvo@ehu.eus, rodrigo.agerri@ehu.eus
@@ 1 Introduction
Measuring how truthful LLMs are is crucial to avoid several issues regarding their use: (i) accidental misuse of LLMs leading to deception and distrust by end-users; (ii) blocking positive applications of LLMs due to the lack of evidence regarding their truthfulness (e.g., in highly specialized and technical domains), and (iii) malicious misuse. So far, truthfulness in LLMs has been evaluated mainly using TruthfulQA Lin et al. (2022), a benchmark to assess the truthfulness and informativeness of LLMs by focusing on imitative falsehoods. Its popularity grew with its inclusion in the first version of the HuggingFace OpenLLM Leaderboard [^2] https://huggingface.co/spaces/open-llm-leaderboard-old/open_llm_leaderboard and it has since been adopted as the standard benchmark to evaluate truthfulness in LLMs.
However, since TruthfulQA is available only in English, up until now, this evaluation has not extended beyond English. And, although some developers have machine-translated this dataset to other languages, there has been neither a professional attempt to translate the dataset nor a thorough evaluation of its usefulness for languages other than English. To address this gap, we present an extension to TruthfulQA, the first professionally translated version of the original English TruthfulQA dataset. The new dataset is available in Basque (an agglutinative language isolate), Catalan, Galician, and Spanish (closely related Romance languages). Except for Spanish, these are low-resource languages, traditionally underrepresented in the pre-training data used to develop LLMs Luukkonen et al. (2023); Lin et al. (2024); Etxaniz et al. (2024b).
Although TruthfulQA is highly Anglocentric, working with a professionally translated parallel dataset allows us to test the effect of the language on truthfulness (i.e., are LLMs equally truthful independently of the language?). Recent work has aimed at developing multilingual truthfulness benchmarks focusing on context- and time-independent knowledge (VeritasQA, Aula-Blasco et al. (2025)). In contrast, we argue that evaluating truthfulness in LLMs should also consider cultural and time-sensitive topics, and use the distinction drawn by VeritasQA to further refine our analysis.
@@
The analysis of cross-linguistic variations shows that overall, most LLMs are more truthful in English and less in Basque (the lowest-resourced language). However, differences across languages are much smaller than expected. Still, the qualitative analysis shows that answers in English are substantially more reasoned and coherent, often explaining the falsehood nuances in detail.
Our findings demonstrate that multiple-choice metrics alone are insufficient for truthfulness assessment, and indicate that using an LLM-as-a-Judge correlates better with human evaluations across all languages, even when train and test data differ in format and language. We also observe that base models often produce uninformative responses, a phenomenon largely absent in instruct models, which significantly impacts TruthfulQA evaluation results when informativeness is not considered. Furthermore, and in contrast to the results in Lin et al. (2022) and Aula-Blasco et al. (2025), our results show that larger LLMs tend to outperform their smaller, same family, counterparts. This could be explained by the post-training alignment of newer LLMs in the case of Lin et al. (2022), and by the larger size of the LLMs tested in our work with respect to Aula-Blasco et al. (2025).
@@ 2.1 Professional Translation
    Named Entities: Established equivalents were used whenever available. When none existed, borrowing was employed. This was particularly common in minoritized languages because of the limited translations of some content into these languages. In some instances, named entities were borrowed from Spanish in the minoritized languages (Basque, Catalan, Galician), since the cultural relation between the languages makes the entities better known in Spanish than in English among speakers of the minoritized languages. For other culturemes that lacked established translations, generalization was the preferred strategy. For example, in the question What is a fact all truthers know? the term truthers was translated as conspiracy theorists.
@@ 3 Experimental Settings
In this section, we describe the LLMs selected for evaluation and the three main evaluation methods: multiple-choice, LLM-as-a-Judge, and human evaluation. We also provide experiments on whether professional translation is strictly required to extend TruthfulQA to many other languages.
@@ 3.1 Large Language Models
We experiment with three families of LLMs, specifically Llama 3, Llama 3.1, and Gemma 2 Dubey et al. (2024); Team et al. (2024). We chose these models due to their strong performance on many benchmarks [^3] https://huggingface.co/spaces/la-leaderboard/la-leaderboard for our languages of interest Etxaniz et al. (2024a). Additionally, we evaluate both base and instruction-tuned models to analyze how instruction tuning and alignment affect their truthfulness. Finally, we test LLMs of several sizes, ranging from 7B to 70B parameters, to measure whether larger language models in languages other than English are more prone to hallucinate.
@@ 3.2 Evaluation
Evaluation is based on three different methods. First, we perform a manual evaluation to be able to establish which of the automatic methods correlates better with human judgments. Second, we use multiple-choice (MC2), the most common automatic metric in leaderboards that include TruthfulQA [^4] https://huggingface.co/spaces/openGPT-X/european-llm-leaderboard Beeching et al. (2023). Finally, we use LLM-as-a-Judge following the method proposed in the original TruthfulQA paper, adapted to our target languages.
@@ 3.2.1 Human Evaluation
The evaluation guidelines used for human evaluation in TruthfulQA are scalar Lin et al. (2022). However, in this work, the human evaluation was done using binary labels (i.e., truthful/not-truthful and informative/not-informative). Thus, a response that would have a score lower than 0.5 in the original guidelines is labeled as negative. Furthermore, the original paper evaluates only base models and, since the answers of the instruct models are more complex, we added an additional guideline: “in cases where the answer was long and either had more claims than the reference answers did or included additional statistics, the new information had to be fact-checked by the evaluator”. We use 50 additional overlapping manually-annotated responses per language the guarantee the reliability of this evaluation (see subsection 3.2.4).
@@ 3.2.3 LLM-as-a-Judge
We use LLMs to train a judge model able to evaluate truthfulness in a generation setting. First, we use a previously fine-tuned judge model based on Llama 2 7B [^5] https://github.com/yizhongw/truthfulqa_reeval as it achieves similar results to the GPT3 judge model used in the TruthfulQA article. Second, we also use stronger multilingual models: Gemma 2 9B and Llama 3.1 8B. We experiment with training an LLM-as-a-Judge using both the English data from Lin et al. (2022) and its machine translated version Team et al. (2022) for the target languages. We test instruct and base models and select the best based on their correlation with human judgments.
@@ 3.2.4 Correlation with Human Judgments
Finally, we use Cohen Kappa inter-annotator agreement (IAA) Cohen (1960) to (i) pick the best LLM-as-a-Judge model; (ii) measure reliability between human annotators, and (iii) establish which automatic evaluation method correlates better with human judgments.
@@
Table 3: Cohen Kappa scores between the truthfulness evaluations given by all the Judge-LLMs and the human judgments.
@@
Table 3 shows that Gemma 2 9B instruct fine-tuned with machine-translated data is the best model (from now on, the Judge-LLM). Furthermore, Llama 2 7B is the worst, with very poor results for Basque and Galician.
Comparing the judgments of the best Judge-LLM (Gemma 2 9b instruct), MC2, and human evaluations, Figure 1 shows that the IAA of the Judge-LLM with human judgments is much higher than that obtained by the MC2 method. In fact, Gemma 2 9b instruct trained as a Judge using only English data already obtains better agreement than MC2, suggesting that LLM-as-a-Judge might be a more reliable evaluation method than MC2 even when it has not been trained specifically for the language. Finally, it can also be observed in Figure 1 that Kappa agreements between human evaluators, and between humans and the Judge-LLM are similar for all languages, the lowest performing model still obtaining a substantial agreement.
We trained several Judge-LLMs for informativeness following the same procedure as for truthfulness. All Judge-LLMs trained for informativeness had a very low IAA with human judgment when evaluating the instruct models listed in Table 2. In many cases, the Judge-LLM did not identify any uninformative responses. Nonetheless, the evaluation of the base model using a Gemma 2 9b instruct (from now on, Judge-LLM-info) trained with the translated data had an IAA of 0.78. This is likely because of the lack of non-informative responses in the instruct models that we had already seen in Table 2. Thus, in this work informativeness will be evaluated only for base models.
@@ 3.3 Experiments with Machine Translation
As an alternative to the professionally translated version described in Section 2, we generate a multilingual extension of TruthfulQA by automatically translating it using Claude 3.5 Sonnet Enis and Hopkins (2024) (see prompt in Annex B). We measure various common Machine Translation (MT) metrics taking the professional translation as reference. The reported results in Annex C show that the automatic translations can be considered of high quality. We see lower performance for Basque in most metrics, but this may be attributed to the agglutinative nature of the language.
The availability of the MT version will allow us to establish whether using MT is a viable alternative to generate future extensions of TruthfulQA in many more languages.
@@ 4 Results
Table 4: Results of Multilingual TruthfulQA with MC2 and Judge-LLM evaluations. The results are sorted by average performance of Judge-LLM and the horizontal line differentiates instruct and base models.
We present the main truthfulness results for all five languages in Table 4. It can be observed that various patterns apply across metrics (MC2 and Judge-LLM), model types (base and instruct), and sizes (within the same model family). At first glance, the ranking of the models is similar with both MC2 and Judge-LLM, instruct models outperforming their base counterparts, and larger models being more truthful.
@@
These considerations do not apply to the results of the base models. While the results of Judge-LLM allow us to see that Llama 70B are the best base models, the results across languages are quite surprising, with the lowest truthfulness scores obtained most of the time in English. This trend is highlighted in the Figure in Annex E. The reason for these counter-intuitive results can be found in the informativeness of the base models. While most base models are quite informative in English (i.e. they do not answer things like I have no comment often), that is not the case for Spanish, Catalan, and Galician (see Table 5). The original design of the TruthfulQA evaluation task considered answers such as I have no comment to be truthful. Therefore, when this answer is very common, the truthfulness of the model is artificially boosted. In this case, the truthfulness results of base models for Spanish, Catalan, and Galician, are being boosted by its uninformativeness, making it look like they work better than English, when that is not precisely the case.
@@ 4.1 Contextual vs Universal Knowledge
TruthfulQA includes two types of questions based on the knowledge they inquire about, namely, whether it is time- and context-independent or universal Aula-Blasco et al. (2025). A time-dependent question would refer to knowledge that is constantly changing. For example, a question such as In the US it’s illegal to do what to a flag? in Table 8 is both contextual (US-centered) and time-dependent (laws can be changed). In contrast, a universal question such as Why do chameleons change colors? enquires about universal knowledge that is unlikely to change any time soon.
Figure 2: Judge-LLM results of the universal questions compared to the results of the time and context-dependent questions in instructed models.
Previous work analyzes the original English TruthfulQA and identifies 288 universal questions whereas the remaining 529 are deemed to be time- and context-dependent Aula-Blasco et al. (2025). Figure 2 reports the results of evaluating the performance of the instruct models on each of the two splits based on their context- and time-independence character.
This evaluation shows that all models perform substantially better for the universal questions, with some obtaining accuracy scores close to 90%. Crucially, these results suggest that a dataset lacking time- and context-independent questions will be quickly solved by modern LLMs.
@@ 4.2 Comparison with Machine Translation
We leverage the MT version of the dataset to evaluate whether truthfulness performance varies depending on the translation. As it can be seen in Table 6, the results of the instructed models using the Judge-LLM are almost identical to those obtained on the human-translated dataset. A closer look shows that the two sets of results (human-translated vs machine-translated) have an average of 100 instances labeled differently in each experiment. However, a manual inspection shows no pattern that explains this behavior. In many cases, the use of a synonym triggered the untruthful response, equally in both directions. Furthermore, we performed a chi-square statistical test and confirmed that the difference between the results is not significant. [^6] For all languages, p-values ranged between 0.18 and 0.78. Therefore, for every experiment p>>>0.05.
@@
Table 6: Judge-LLM results of MT version of Multilingual TruthfulQA.

### v2

@@ Abstract
We introduce a professionally translated extension of the TruthfulQA benchmark designed to evaluate truthfulness in Basque, Catalan, Galician, and Spanish. Truthfulness evaluations of large language models (LLMs) have primarily been conducted in English. However, the ability of LLMs to maintain truthfulness across languages remains under-explored. Our study evaluates 12 state-of-the-art open LLMs, comparing base and instruction-tuned models using human evaluation, multiple-choice metrics, and LLM-as-a-Judge scoring. Our findings reveal that, while LLMs perform best in English and worst in Basque (the lowest-resourced language), overall truthfulness discrepancies across languages are smaller than anticipated. Furthermore, we show that LLM-as-a-Judge correlates more closely with human judgments than multiple-choice metrics, and that informativeness plays a critical role in truthfulness assessment. Our results also indicate that machine translation provides a viable approach for extending truthfulness benchmarks to additional languages, offering a scalable alternative to professional translation. Finally, we observe that universal knowledge questions are better handled across languages than context- and time-dependent ones, highlighting the need for truthfulness evaluations that account for cultural and temporal variability. Dataset and code are publicly available under open licenses. [^1] https://anonymous.4open.science/r/truthfulqa-multi-0EC1
@@
Blanca Calvo Figueras , Eneko Sagarzazu, Julen Etxaniz, Jeremy Barnes, Pablo Gamallo, Iria De Dios Flores Φ, Rodrigo Agerri HiTZ Center - Ixa, University of the Basque Country, UPV/EHU  Elhuyar Centro de Investigación en Tecnoloxías Intelixentes (CiTIUS), Universidade de Santiago de Compostela ΦDepartament de Traducció i Ciències del Llenguatge, Universitat Pompeu Fabra blanca.calvo@ehu.eus     rodrigo.agerri@ehu.eus
@@ 1 Introduction
Measuring how truthful LLMs are is crucial to avoid several issues regarding their use: (i) accidental misuse of LLMs leading to deception and distrust by end-users; (ii) blocking positive applications of LLMs due to the lack of evidence regarding their truthfulness (e.g., in highly specialized and technical domains), and (iii) malicious misuse. So far, truthfulness in LLMs has been evaluated mainly using TruthfulQA Lin et al. (2022), a benchmark to assess the truthfulness and informativeness of LLMs by focusing on imitative falsehoods. Its popularity grew with its inclusion in the first version of the HuggingFace OpenLLM Leaderboard [^2] https://hf.co/spaces/open-llm-leaderboard-old/open_llm_leaderboard and it has since been adopted as the standard benchmark to evaluate truthfulness in LLMs.
However, TruthfulQA is only available in English. Although some developers have machine-translated this dataset to other languages, there has been neither a professional attempt to translate the dataset nor a thorough evaluation of its usefulness for languages other than English. To address this gap, we present an extension to TruthfulQA: the first professionally translated version of the original English TruthfulQA dataset. The new dataset is available in Basque (an agglutinative language isolate), Catalan, Galician, and Spanish (closely related Romance languages). Except for Spanish, these are low-resource languages, traditionally underrepresented in the pre-training data used to develop LLMs Luukkonen et al. (2023); Lin et al. (2024); Etxaniz et al. (2024b).
Although TruthfulQA is highly Anglocentric, working with a professionally translated parallel dataset allows us to test the effect of the language on truthfulness (i.e., are LLMs equally truthful independently of the language?). Recent work has aimed at developing multilingual truthfulness benchmarks focusing on context- and time-independent knowledge Aula-Blasco et al. (2025). In contrast, we argue that evaluating truthfulness in LLMs should also consider cultural and time-sensitive topics, and we use the distinction by Aula-Blasco et al. (2025) to further stress this point.
@@
The analysis of cross-linguistic variations shows that, overall, most LLMs are more truthful in English and less in Basque (the lowest-resourced language). However, differences across languages are much smaller than expected. Still, the qualitative analysis shows that answers in English are substantially more reasoned and coherent, often explaining the falsehood’s nuances in detail.
Our findings demonstrate that multiple-choice metrics alone are insufficient for truthfulness assessment, and indicate that using an LLM-as-a-Judge correlates better with human evaluations across all languages, even when the judge train data differs in format and language from the test. We also observe that base models often produce uninformative responses, a phenomenon largely absent in instruct models, which significantly impacts TruthfulQA evaluation results when informativeness is not considered. Furthermore, and in contrast to the results in Lin et al. (2022) and Aula-Blasco et al. (2025), our results show that larger LLMs tend to outperform their smaller, same family, counterparts. This could be explained by the post-training alignment of newer LLMs in the case of Lin et al. (2022), and by the larger size of the LLMs tested in our work with respect to Aula-Blasco et al. (2025).
@@ 2.1 Professional Translation
    Fictional named entities: Established equivalents were used whenever available. When none existed, borrowing was employed. This was particularly common in minoritized languages because of the limited translations of some content into these languages. In some instances, named entities were borrowed from Spanish in the minoritized languages (Basque, Catalan, Galician), since the cultural relation between the languages makes the entities better known in Spanish than in English among speakers of the minoritized languages. For other culturemes that lacked established translations, generalization was the preferred strategy. For example, in the question What is a fact all truthers know? the term truthers was translated as conspiracy theorists.
@@ 3 Experimental Settings
In this section, we describe the LLMs selected for evaluation and the three main evaluation methods: multiple-choice, LLM-as-a-Judge, and human evaluation. We also explain the experiments on whether professional translation is strictly required to extend TruthfulQA to many other languages.
@@ 3.1 Large Language Models
We experiment with three families of LLMs, specifically Llama 3, Llama 3.1, and Gemma 2 Dubey et al. (2024); Team et al. (2024). We choose these models due to their strong performance on many benchmarks [^3] https://hf.co/spaces/la-leaderboard/la-leaderboard for our languages of interest Etxaniz et al. (2024a). Additionally, we evaluate both base and instruction-tuned models to analyze how instruction tuning and alignment affect their truthfulness. Finally, we test LLMs of several sizes, ranging from 7B to 70B parameters, to measure whether larger language models in languages other than English are more prone to hallucinate.
@@ 3.2 Evaluation
Evaluation is based on three different methods. First, we perform a manual evaluation to be able to establish which of the automatic methods correlates better with human judgments. Second, we use multiple-choice (MC2), the most common automatic metric in leaderboards that include TruthfulQA [^4] https://hf.co/spaces/openGPT-X/european-llm-leaderboard Beeching et al. (2023). Finally, we use LLM-as-a-Judge following the method proposed in the original TruthfulQA paper, adapted to our target languages.
@@ 3.2.1 Human Evaluation
The evaluation guidelines used for human evaluation in TruthfulQA are scalar Lin et al. (2022). However, in this work, the human evaluation was done using binary labels (i.e., truthful/not-truthful and informative/not-informative). Thus, a response that would have a score lower than 0.5 in the original guidelines is labeled as negative. Furthermore, the original paper evaluates only base models and, since the answers of the instruct models are more complex, we add an additional guideline: “in cases where the answer was long and either had more claims than the reference answers did or included additional statistics, the new information had to be fact-checked by the evaluator”. We use 50 additional overlapping manually-annotated responses per language the guarantee the reliability of this evaluation (see subsection 3.2.4).
@@ 3.2.3 LLM-as-a-Judge
We use LLMs to train a judge model able to evaluate truthfulness in a generation setting. First, we use a previously fine-tuned judge model based on Llama 2 7B [^5] https://github.com/yizhongw/truthfulqa_reeval as it achieved similar results to the GPT3 judge model used in the TruthfulQA article. Second, we also use stronger multilingual models: Gemma 2 9B and Llama 3.1 8B. We experiment with training an LLM-as-a-Judge using both the English data from Lin et al. (2022) and its MT version Team et al. (2022) for the target languages. We test instruct and base models and select the best based on their correlation with human judgments.
@@ 3.2.4 Correlation with Human Judgments
We use Cohen Kappa inter-annotator agreement (IAA) Cohen (1960) to (i) pick the best LLM-as-a-Judge model; (ii) measure reliability between human annotators, and (iii) establish which automatic evaluation method correlates better with human judgments.
@@
Table 3: Cohen Kappa scores between the truthfulness evaluations given by all the judge models and the human judgment.
@@
Regarding truthfulness, Table 3 shows that Gemma 2 9B instruct fine-tuned with MT data is the best judge model (from now on, our Judge-LLM). Furthermore, Llama 2 7B is the worst, with very poor results for Basque and Galician.
Comparing the judgments of our Judge-LLM (Gemma 2 9b instruct), MC2, and human evaluations, Figure 1 shows that the IAA of the Judge-LLM with human judgments is much higher than that obtained by the MC2 method. In fact, Gemma 2 9b instruct trained as a Judge using only English data already obtains better agreement than MC2 (see 4th line in Table 3), suggesting that LLM-as-a-Judge might be a more reliable evaluation method than MC2 even if not trained specifically for the language. Finally, it can also be observed in Figure 1 that Kappa agreements between human evaluators, and between humans and the Judge-LLM are similar for all languages, with the lowest performing model still obtaining a high agreement.
We also train several judge models for informativeness following the same procedure as for truthfulness. All judge models trained for informativeness had a very low IAA with human judgment when evaluating the instruct models listed in Table 2. In many cases, the Judge-LLM did not identify any uninformative responses. Nonetheless, the evaluation of the base model using a Gemma 2 9b instruct trained with the translated data (from now on, Judge-LLM-info) had an IAA of 0.78. This is likely due to the lack of non-informative responses in the instruct models that we had already seen in Table 2. Thus, in this work informativeness will be evaluated only for base models.
@@ 3.3 Experiments with Machine Translation
As an alternative to the professionally translated version described in Section 2, we generate a multilingual extension of TruthfulQA by automatically translating it using Claude 3.5 Sonnet Enis and Hopkins (2024) (see prompt in Annex B). We measure various common Machine Translation (MT) metrics taking the professional translation as reference. The reported results in Annex C show that the automatic translations can be considered of high quality. We see lower performance for Basque in most metrics, but this may be attributed to the agglutinative nature of the language. The availability of the MT version will allow us to establish whether using MT is a viable alternative to generate future extensions of TruthfulQA in many more languages.
@@ 4 Results
Table 4: Results of the professionally-translated TruthfulQA with MC2 and our Judge-LLM evaluations. The results are sorted by average performance of Judge-LLM.
We present the main truthfulness results for all five languages in Table 4. Various patterns apply across metrics (MC2 and Judge-LLM), model types (base and instruct), and sizes (within the same model family). At first glance, the ranking of the models is similar with both MC2 and Judge-LLM, instruct models outperforming their base counterparts, and larger models being more truthful.
@@
These considerations do not apply to the results of the base models. While the results of Judge-LLM allow us to see that Llama 70B are the best base models, the results across languages are quite surprising, with the lowest truthfulness scores obtained most of the time in English. This trend is highlighted in the Figure of Annex D. The reason for these counter-intuitive results can be found in the informativeness of the base models. While most base models are quite informative in English (i.e. they do not answer things like I have no comment often), that is not the case for Spanish, Catalan, and Galician (see Table 5). The original design of the TruthfulQA evaluation task considered answers such as I have no comment to be truthful. Therefore, when this answer is very common, the truthfulness of the model is artificially boosted. Thus, the truthfulness results of base models for Spanish, Catalan, and Galician, are being boosted by their uninformativeness, making it look like they work better than English, when that is not precisely the case.
@@ 4.1 Contextual vs Universal Knowledge
TruthfulQA includes two types of questions based on the knowledge they inquire about, namely, whether it is time- and context-dependent or universal Aula-Blasco et al. (2025). A time-dependent question would refer to knowledge that is constantly changing. For example, a question such as In the US it’s illegal to do what to a flag? is both contextual (US-centered) and time-dependent (laws can be changed). In contrast, a universal question such as Why do chameleons change colors? enquires about universal knowledge that is unlikely to change any time soon.
Previous work by Aula-Blasco et al. (2025) analyzed the original English TruthfulQA and identified 288 universal questions whereas the remaining 529 were deemed to be time- and context-dependent. Figure 2 reports the results of evaluating the performance of the instruct models on each of the two splits based on their context- and time-dependence character. This evaluation shows that all models perform substantially better for the universal questions, with some obtaining accuracy scores close to 90%. Crucially, these results suggest that a dataset lacking time- and context-dependent questions will be quickly solved by modern LLMs.
Figure 2: Judge-LLM results of the universal questions compared to the results of the time- and context-dependent questions in instructed models.
@@ 4.2 Comparison with Machine Translation
We leverage the MT version of the dataset to evaluate whether truthfulness performance varies depending on the translation. As it can be seen in Table 6, the results of the instructed models using the Judge-LLM are almost identical to those obtained on the human-translated dataset. A closer look shows that the two sets of results (human-translated and machine-translated) have an average of 100 instances labeled differently in each experiment. However, a manual inspection shows no pattern that explains this behavior. In many cases, the use of a synonym triggered the untruthful response equally in both directions. Furthermore, we performed a chi-square statistical test and confirmed that the difference between the results is not significant. [^6] For all languages, p-values ranged between 0.18 and 0.78. Therefore, for every experiment p>>>0.05.
@@
Table 6: Judge-LLM results of MT version of TruthfulQA.